OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6). 12 13 -11 -7

OpenStudy (unklerhaukus):

\[f(x)= -2x + 1\] \[f(-6)=-2(-6)+1\]\[~~~~~~=12+1\]\[~~~~~=13\]